OpenStudy (anonymous):

how does 10 log of base 10(10^5+ 10^6) become 10 log of base 10 (1.1 * 10^6) ?

OpenStudy (unklerhaukus):

10^5 = 100 000 10^6 = 1 000 000 10^5+10^6 = 1 100 000 = 1.1x10^6

OpenStudy (anonymous):

10^5 + 10^6 = 10^5 (1+10) = 10^5 (11) = (1.1) 10^6